OpenStudy (anonymous):

Help me; I don't know where to start. Prove: 1) cotx- (cos2x/sinx cosx) = tanx 2) Sinx/sin^2x+cos2x = secx/cotx 3) (cscx(cosx+sinx)^2)/1+sin2x = cotx/cscx

OpenStudy (anonymous):

Let's rewrite the equation: cotx - tanx = (cos2x/sinxcosx) Let's look at (cos2x/sinxcosx) Use the double angle formula: \[\cos2x = \cos ^{2}x - \sin ^{2}x\] Split up the numerators while keeping a common denominator: \[[(\cos ^{2}x)/(sinxcosx)] - (\sin ^{2}x)/(sinxcosx)]\] After canceling we are left with: (cosx/sinx) - (sinx/cosx) Use the reciprocal identities: cosx/sinx = cotx and sinx/cosx = tanx So, this reduces to cotx-tanx

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

oh...wait. I get it. thank you.

OpenStudy (anonymous):

If you need help with the other two, let me know

OpenStudy (anonymous):

are we allowed to rewrite the equation to cotx - tanx = (cos2x/sinxcosx) ?

OpenStudy (anonymous):

yes, of course!

OpenStudy (anonymous):

oh really? I was never told that... o_o

OpenStudy (anonymous):

Do you think you could help me with (cscx(cosx+sinx)^2)/1+sin2x = cotx/cscx I think I got the 2nd problem but this one's confusing

OpenStudy (anonymous):

simplify (cosx=sinx)^2 first. The rest will follow.

OpenStudy (anonymous):

should read (cosx+sinx)^2

OpenStudy (anonymous):

so it'll be cos^2x+2cosxsinx+sin^2x?

OpenStudy (anonymous):

Yes. now what is cos^2x+sin^2x ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so what do you have?

OpenStudy (anonymous):

1+2cosxsinx

OpenStudy (anonymous):

Next what is sin2x ?

OpenStudy (anonymous):

1-cos^2x

OpenStudy (anonymous):

wait opps...nvm

OpenStudy (anonymous):

it's 2sinxcosx

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Can you simplify the whole equation now?

OpenStudy (anonymous):

so do they cross out and leave cscx? or....

OpenStudy (anonymous):

because I dont think they'll cross out since cscx is multiplying it

OpenStudy (anonymous):

just write down the equation first

OpenStudy (anonymous):

cscx(1+2sinxcosx)/1+2sinxcosx --> cscx+2sinxcosxcscx/1+2sinxcosx?

OpenStudy (anonymous):

they will cancel out. what you will have left is cscx on the left hand side

OpenStudy (anonymous):

Now simplify secx/cotx

OpenStudy (anonymous):

Remember secx=(1/cosx) and cotx = (cosx)/(sinx)

OpenStudy (anonymous):

where did you get the sec from?

OpenStudy (anonymous):

from the right hand side of your original equation

OpenStudy (anonymous):

there's no sec

OpenStudy (anonymous):

sorry, i was looking at #2

OpenStudy (anonymous):

it's cotx/cscx on the right side

OpenStudy (anonymous):

cotx/cscx

OpenStudy (anonymous):

so how does cscx alone get to cotx/cscx?

OpenStudy (anonymous):

hmm! let me take a look again

OpenStudy (anonymous):

give me a few minutes

OpenStudy (anonymous):

I gotta go but try simplifying the equations again. We might have missed something.